(square root of 2x + 1) +1 = 7
Answer by unlockmath(1688)

Hello,
Here's what we do with this: (square root of 2x + 1) +1 = 7
Subtract 1 from both sides to get:
(square root of 2x + 1) = 6
Now square both sides to get:
2x+1=36
Subtract 1 and divide 2 on each side:
x=35/2 or 14 1/2
Plug it in and we see it's correct.
Make sense?
RJ
www.math-unlock.com